Fermi National Laboratory
 
Operation Summary for January 9, 2012 - January 11, 2012
 

Monday January 9

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, a Water Works person reported at 8:22 AM that he was isolating the MI LCW system from CUB to regenerate column #1. Operations turned off the Booster beam and began preparing the area for access at 9:15 AM. Operators loaded an Energy Conservation
Timeline into the Time Line Generator (TLG) at 9:18 AM. Booster RF personnel began working on an RF station (BRF18) around 9:45 AM. Linac personnel began a study. The Booster RF worked ended and Operations had the Booster back on by 11:52 AM. Operations had established beam to the NuMI and MiniBooNE experiments, and to the Meson FTBF experiment T-958 by 12:18 PM.

During the evening shift, T-958 made a number of accesses throughout the shift as they made changes to their experiment. Operations disabled their beam at midnight. Operators tuned the Booster near the end of shift.

Tuesday January 10

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 1:57 AM, Operators reported that the MI-60 magnet LCW system was losing water. Since the MI system at CUB was put into regeneration mode on Monday, there was no way to make up water to the system. Operations called in many people to search for the leak. Operators found no leaks in the Service Buildings or at CUB. Operators turned off the MI LCW pumps at 3:11 AM and switched off MI for access at 3:23 AM. A Water Group person reported finding the leak on a hose in MI-30 tunnel near the collimators. At 4:41 AM, Operators reported finding a small leak on a quadrupole magnet (Q861) in the MiniBooNE line. Water Group personnel repaired all the leaks. At 7:43 AM, Operators reported that the LCW pumps at MI-20 and MI-30 would not come back on. Operations established beam to the Booster dump at 7:55 AM.

During the day shift, a Water Group person reported that the problems with the LCW pumps at MI-20 and MI-30 would take several hours to diagnose and repair. Water Group personnel reported at 8:06 AM that the CUB column #1 was back online and they were making up water to the MI LCW system. Operations turned on MI and had the accelerator ramping by 8:39 AM. EE Support personnel began checking MI-60 north and south at 9:04 AM for LCW leaks that might have occurred when the LCW pumps were turned back on; they didn’t find any. Operations had beam established to NuMI and MiniBooNE by 9:24 AM. Operators loaded protons into the Recycler at 9:39 AM. Booster personnel began a two-hour aperture scan study at 9:42 AM. Operations established beam to T-958 at 11:06 AM. Muon Rings personnel began a study period at 12:30 PM. Operations received a hot magnet indication for a quadrupole magnet (Q310) in MI at 12:46 PM; Operators prepared the area for an access. Water Group personnel reported at 1:48 PM that they found a closed LCW valve in the tunnel. Operations turned on MI and had reestablished beam to MiniBooNE and NuMI by 2:34 PM.

During the evening shift, the Muon Rings study ended at 4:58 PM. Operators loaded protons into the Recycler. Operations received permission to send beam to the next Meson FTBF experiment T-992, which should began taking beam on Wednesday (1/11/12) morning. At 9:14 PM, a Booster vacuum pump (B:IPRF15) tripped off and took down two Booster RF stations (BRF3 and BRF4); everything reset. Operations disabled T-958 beam at 10:04 PM; this ended their current run.

Wednesday January 11

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, Operations completed a tune of Linac at 12:30 AM. At 2:08 AM, the NuMI beam permit tripped off on a MI incomplete RF gap indication; it reset. Interlock personnel reported at 7:21 AM that they would be conducting safety system tests for Switchyard.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo approved of the plans by Booster and Muon Rings personnel to conduct studies. The OpsCo also reported that on 1/18/12, there would be an eight-hour access in the Booster tunnel for solid-state modulator and power amplifier upgrades.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/11/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory